Grade 11Algebra

If the am and hm of two number of 4 and 3 respectively then find the number

Smruti Ludu
7 Years agoGrade 11

1 Answer

Baijnath
7 Years ago
Given HM=2ab/(a+b)=3..............(1)
AM=(a+b)/2=4......(2)
=a+b=8.......(3)
Putting this in eqn  1..we get 
ab=12...............(4)
Now...(a-b)^2=(a+b)^2-4ab
=(a-b)^2=(8)^2-4×12
=(a-b)^2=16
Hence,(a-b)=4......(5)
Solving 3 and 5...we get
a=6 &b=2
Viz your required answer!!!
Hope u understand😊😊😊😊
Regars ,Baijnath